### v1

@@ Abstract
Context. Stars with initial mass above roughly 8 $M_{\odot}$ will evolve to form a core made of iron group elements at which point no further exothermic nuclear reactions between charged nuclei may prevent the core collapse. Electron captures, neutrino losses, and the photo-disintegration of heavy nuclei trigger the collapse of these stars. Models at the brink of core collapse are produced using stellar evolution codes and these pre-collapse models may be used in the study of the subsequent dynamical evolution (including their explosion as supernovae and the formation of compact remnants such as neutron stars or black holes).
Aims. We upgrade the physical ingredients employed by the GENeva stellar Evolution Code, GENEC, so that it may cover the regime of high temperatures and high densities required to produce progenitors of core-collapse. Our ultimate goal is producing suitable pre-supernova models with GENEC, not only right before collapse, but also during late phases (silicon/oxygen burning).
Methods. We have improved GENEC in three directions, equation of state, the nuclear reaction network and the radiative and conductive opacities adapted for the computation of the advanced phases of evolution. We produce a small grid of pre-supernova models of stars with zero-age main sequence masses of $15M_{\odot}$, $20M_{\odot}$ and $25M_{\odot}$ at solar and less than half solar metallicities. The results are compared with analogous models produced with the MESA code.
Results. The global properties of our new models, particularly of their inner cores, are comparable to models computed with MESA and pre-existing progenitors in the literature. Between codes the exact shell structure varies impacting explosion predictions.
Conclusions. Using GENEC with state-of-the-art physics, we have produced massive stellar progenitors prior to collapse. These progenitors are suitable for follow-up studies, including the dynamical collapse and supernova phases. Larger grids of supernova progenitors are now feasible, with potential for further dynamical evolution.
@@ 1 Introduction
Building stellar evolution models to reproduce observations remains a difficult task due to a relatively large amount of free parameters and approximations used in the modelling. An increase of the quality and quantity of observations of stellar objects, (e.g., the increasing catalog of observations by the GAIA collaboration, Gaia Collaboration et al. 2016, 2018, 2021) refines the global constraints that stellar models must satisfy (most of which refer to photospheric conditions), as can be found in recent studies (Dal Tio et al. 2022; Massey et al. 2023; Yang et al. 2019). Development of asteroseismology techniques, reviewed by Aerts et al. (2019), can provide an insight into the internal characteristics of stars that also helps constrain stellar modelling (e.g., Burssens et al. 2023; Bowman 2023; Eggenberger et al. 2019a, b). Furthermore, the observations at ever larger redshifts provided by the JWST give hints about the early stellar generations in the universe that may have produced, for example, the seeds of supermassive black holes or the strongly nitrogen enriched regions (see for instance Ji et al. 2024; Jeon et al. 2024; Nandal et al. 2024a, b). Among the most prominent tasks for stellar models is to verify the constraints provided from all of these observations.
@@
One major difficulty, for massive stars above 8 $M_{\odot}$, is that the stellar core in the final phases of evolution, i.e., from the beginning of carbon fusion in the core, is obscured from view deep inside the large stellar envelopes. The impact of mixing and nuclear burning in the stellar interior therefore cannot be well constrained by observations. Understanding the evolution of the stellar core in this regime heavily relies on numerical models, whose validation chiefly depends on observations of the subsequent evolution. Namely, pre-supernova models are further employed as the starting points to predict the fate after core bounce and through to the supernova explosions (e.g., Ugliano et al. 2012; Sukhbold & Woosley 2014; Sukhbold et al. 2016, 2018) as well as the multi-messenger signal of the post-explosion evolution (e.g, Murphy et al. 2009; Yakunin et al. 2010; Scheidegger et al. 2010; Mueller et al. 2013; Lund et al. 2012; Cerdá-Durán et al. 2013; Tamborra et al. 2013, 2014b, 2014a; Kuroda et al. 2016; Mirizzi et al. 2016; Kashiyama et al. 2016; Janka et al. 2016; Andresen et al. 2017; Kuroda et al. 2017; Janka 2016; Kuroda et al. 2017; Pan et al. 2018; O’Connor & Couch 2018; Powell & Müller 2019; Radice et al. 2019; Morozova et al. 2018; Walk et al. 2018; Andresen et al. 2019; Lin et al. 2020; Walk et al. 2019; Müller 2019; Andersen et al. 2021; Powell & Müller 2020; Pan et al. 2021; Mezzacappa et al. 2020; Schneider et al. 2021; Andersen et al. 2021; Jardine et al. 2022; Mezzacappa et al. 2023; Bugli et al. 2023, 2023; Vartanyan et al. 2023). With these predictions, we may confront observations of the observed statistics of neutron stars and black holes, including constraints on their spins, supernova explosion properties, predicted gravitational waveforms, neutrinos or nucleosynthetic yields with the obtained results and, hence, constrain the pre-supernova models (e.g., Smartt et al. 2009; Smartt 2015; Radice et al. 2019; Fonseca et al. 2023; KamLAND et al. 2024; Reichert et al. 2021, 2023, 2024; Sieverding et al. 2023).
@@
Late phase stellar evolution models have been studied by numerous authors using a variety of different stellar evolution codes. With GENEC, Hirschi et al. (2004), pushed models beyond silicon-core burning. For the KEPLER stellar evolution code studies such as Heger et al. (2001a); Sukhbold & Woosley (2014); Sukhbold et al. (2018) have produced a large number of pre-supernova models. The development of MESA in Paxton et al. (2011, 2013, 2015, 2018, 2019) has also produced grids of pre-supernova models. All these studies have provided initial models to probe the subsequent dynamical evolution of collapsing massive stars.
@@
The properties of progenitor stars at the moment of collapse are crucial to determine the outcome of the post collapse evolution of massive stars. Among these properties we mention, e.g., the density, rotational and magnetic profiles, the mass of the iron core and its radius (ordinarily quantified in terms of its compactness parameter O’Connor & Ott 2011), or its electron fraction profile (e.g. Baron & Cooperstein 1990). Supernova explosions are very sensitive to the progenitor structure with the entropy, electron fraction, equation of state and neutrino physics all having a strong impact (Cooperstein & Wambach 1984; Baron et al. 1985; Bethe 1990). The stellar profiles of these quantities result from a complex balance of nuclear burning stages and the creation, spatial dimension, and temporal sequence of different convective layers (some of which may mutually interact). Convection in one-dimensional stellar evolution codes is implemented assembling together the Schwarzchild-Ledoux (local) instability criterion (either the Schwarzshild or the Ledoux criterion are used, but not together), and basic prescriptions for diffusion of chemical species and angular momentum (in addition to parametric models for semi-convection, overshooting, etc.). Since the incorporation of these elements varies from one stellar evolution code to another, more often than desirable, the pre-supernova state for seemingly similar ZAMS conditions differs among different groups (Jones et al. 2015; Chieffi & Limongi 2020).
@@
In this paper, we will show the details of the implementation of an extended equation of state (EoS) in GENEC covering the whole density-temperature plane required to reach core collapse for massive stars. This EoS patches together two pre-existing EoSs, which include complementary physical aspects, along with some common elements (Sect. 2). We will show that while the local values of the EoS recovered thermodynamic variables do not show large discrepancies, their derivatives may have significant deviations, specifically derivatives impacting the criteria used to distinguish among zones unstable or stable to convection. Additionally, we extend the coverage of the radiative and conductive opacities to high temperatures and densities to include the effects of degeneracy and high energy physics, which play an important role in the final phases of stellar life (Sect. 3). With the aim of approximately covering the most advanced burning phases of the evolution of massive stars, we also extend the existing reduced nuclear reaction network to include electron captures on heavy nuclei (Sect. 4). Therefore, the nuclear network approximately accounts for the reduction of the electron fraction ($Y_{e}$) and consequently the electron pressure, which induces core collapse. To validate these improvements, using GENEC, we evolve several models of various massive stars until the brink of core collapse, which we compare to corresponding models evolved with MESA (Sect. 5). We discuss our results and write our conclusions in Sect. 6. Since the implementation details of the previous default EoS in GENEC are not easily accessible in electronic form, we will highlight the most relevant features of that EoS in the Appendix A.
@@
2 A new EoS to cover the regime from partial ionization to relativistic electron degeneration
@@
Stellar matter is usually modelled as a mixture of ideal (baryonic) gases, radiation, and electrons/positrons. It encompasses regimes of partial to full ionisation. The current version of the GENEC stellar evolution code employs, by default, the Dichte [^1] The name Dichte comes from the German word for density as the code was developed within the group of Kippenhahn and the EoS returns the density when one inputs pressure, temperature, and composition. EoS (D-EoS, hereafter), which accounts for the partial ionisation regime of certain light elements, (hydrogen, helium, carbon, oxygen, nitrogen, and magnesium) in the stellar envelope (see Appendix A for details), as well as fully ionised medium of the stellar interior. We note that extending the treatment of partial ionisation for heavier elements would be welcome, as discussed in Jermyn et al. (2021), but this is a matter for future work. Two important regimes for massive star evolution are not covered by the D-EoS, namely the electron degeneracy in the high density (i.e., relativistic) regime and the effects of electron-positron pair creation in the very high temperature regime. To cover this region of the density-temperature parameter space, we have extended the existing GENEC EoS by matching it with the Timmes EoS (Timmes et al. 2000) in the deep stellar interiors. The latter EoS is based upon an interpolation table of the Helmholtz free energy (ensuring thermodynamic consistency) and includes the treatment of positrons. It should be noted that the Timmes EoS (T-EoS hereafter) assumes total ionisation of the plasma, which is not the case in the whole stellar interior. To allow for the EoS to deal with the partial ionisation in the stellar envelope, we employ a switch between the two EoSs. According to Jermyn et al. (2021) the ionisation effects become relevant for temperatures $T<T^{\text{ion}}\approx 10^{4}Z_{j}^{2}\,$ and densities $\rho<\rho^{\text{ion}}\approx 3A_{j}Z_{j}^{3}\,\text{g\,cm}^{-3}$, where $Z_{j}$ and $A_{j}$ are the atomic number and atomic mass of the considered element. Using these criteria on the ionisation conditions as approximate guidelines, and trying to minimize the jump between the two different EoSs, if the following condition is verified, we use the D-EoS,
$\hskip 65.44142pt\rm log_{10}(\rho)\leq 2.8\ \ \rm{or}\ \ \rm log_{10}(T)\leq 7.55,$ (1)
@@
Thus, noting that the switch to full ionization is handled according to Eq. (42) in the D-EoS, the partial ionization regime for hydrogen and helium is entirely addressed by this EoS (and not by the T-EoS). In the following, we will collectively refer to the coupled EoS employed in GENEC as the DT-EoS.
@@ 2.1 Electronic contribution to the pressure
The T-EoS represents an improvement over the D-EoS when it comes to computing the electronic pressure (see Appendix A). While the D-EoS employs analytic estimations for the evaluation of the Fermi-Dirac integrals that are used to compute the electron contribution to the thermodynamic variables, the T-EoS uses a more accurate table for the leptonic contribution including electrons and positrons.
In Fig. 1 we show the variations of the electron-positron pressure contribution as a function of $\rho$, and $T$, between the two EoSs. For the computation of the pressure we take the case of pure $\rm{}^{16}O$, for which $\overline{A}=16$ and $\overline{Z}=8$, although changing the composition does not alter the conclusions made here. The quantity represented in Fig. 1 is,
$\chi=\frac{P_{\rm e}^{T}}{P^{T}}\times\frac{\left|P_{\rm e}^{D}-P_{\rm e}^{T}\right|}{P_{\rm e}^{T}}.$ (4)
@@ 2.2 Impact on density
The change of EoS also impacts the density profile in the star, a crucial element for the determination of the stellar structure. The T-EoS takes as input a pair of values $(\rho,T)$ to construct the Helmholtz free energy and evaluate the pressure $P^{\rm helm}$ (Timmes & Swesty 2000). On the other hand, the D-EoS takes as input the pair of values $(P,T)$ and iterates over the degeneracy parameter, $\Psi$, or electron chemical potential, to find the density ($\rho^{\rm D}$) that satisfies the equations shown in the Appendix A. In Fig. 2 we show the logarithm of the difference $\rho^{\rm D}/\rho^{\rm T}-1$. Where $\rho^{\rm T}$ is the input density for the T-EoS and $\rho^{\rm D}$ the output density of the D-EoS. $\rho^{\rm D}$ is obtained by using as input value $P^{\rm T}$, i.e., the pressure computed using the T-EoS for the thermodynamic conditions $(\rho^{\rm T},T)$. Note that the temperature, $T$, has been fixed in all the process of comparison.
@@
We see in Fig. 2 that in a large part of the $(\rho,T)$ plane $\log_{10}(\rho^{\rm D}/\rho^{\rm T}-1)<10^{-2}$. It is only in the upper left region, where the pressure is dominated mainly by the radiation contribution, that the D-EoS begins to diverge strongly from the thermodynamic consistent T-EoS. Indeed, in this region, the contributions made by the positrons to the pressure can be large. In general, the presence of positrons in the EoS will affect the electron chemical potential and thus the degeneracy parameter. The lack of positrons in the D-EoS, and the analytic expressions used to evaluate the Fermi-Dirac integrals (Appendix A), leads to large differences in the estimation of the degeneracy parameter between the two EoS’. For example, at the point $(\rho,T)=(10^{5}\ \rm g\,cm^{-3},10^{9.5}\ \rm K)$, for a pure silicon composition, the T-EoS returns $\psi^{T}=-1.8$ whereas the D-EoS returns $\psi^{D}=-1.2$. This degeneracy parameter with the D-EoS gives a density of roughly $6\times 10^{5}\,\rm g\,cm^{-3}$ a considerable change to the Timmes input of $1\times 10^{5}\,\rm g\,cm^{-3}.$ Thus using the Timmes EoS in this region is mandatory to obtain a correct evaluation of the density and, thereby, the correct stellar structure.
@@
Most of the stellar interior of our models are not affected by this region of strong differences in the density. However, we overlap in Fig. 2 the $(\rho,T)$ internal structure of a $15M_{\odot}$ model computed with the D-EoS in magenta and using the DT-EoS with a black dashed line at roughly the same point of evolution at the onset of Si core burning. The main shift between these models comes from the location of the interfaces, represented by sharp variations in the $(\rho,T)$ curve. The size of jump at these interfaces in density and temperature is also different between the two models using different EoS’. The properties at shell interfaces are as we see in Sec. 6 crucial for understanding supernova explosions of these progenitors.

### v2

@@ Abstract
Context. Stars with initial mass above roughly 8 $M_{\odot}$ will evolve to form a core made of iron group elements, at which point no further exothermic nuclear reactions between charged nuclei may prevent the core collapse. Electron capture, neutrino losses, and the photo-disintegration of heavy nuclei trigger the collapse of these stars. Models at the brink of core collapse are produced using stellar evolution codes, and these pre-collapse models may be used in the study of the subsequent dynamical evolution (including their explosion as supernovae and the formation of compact remnants such as neutron stars or black holes).
Aims. We upgraded the physical ingredients employed by the GENeva stellar Evolution Code, GENEC, so that it covers the regime of high-temperatures and high-densities required to produce the progenitors of core-collapse. Our ultimate goal is producing pre-supernova models with GENEC, not only right before collapse, but also during the late phases (silicon and oxygen burning).
Methods. We have improved GENEC in three directions: equation of state, the nuclear reaction network, and the radiative and conductive opacities adapted for the computation of the advanced phases of evolution. We produce a small grid of pre-supernova models of stars with zero age main sequence masses of $15M_{\odot}$, $20M_{\odot}$, and $25M_{\odot}$ at solar and less than half solar metallicities. The results are compared with analogous models produced with the MESA code.
Results. The global properties of our new models, particularly of their inner cores, are comparable to models computed with MESA and pre-existing progenitors in the literature. Between codes the exact shell structure varies, and impacts explosion predictions.
Conclusions. Using GENEC with state-of-the-art physics, we have produced massive stellar progenitors prior to collapse. These progenitors are suitable for follow-up studies, including the dynamical collapse and supernova phases. Larger grids of supernova progenitors are now feasible, with the potential for further dynamical evolution.
Key Words.:
Stars: evolution - Stars: massive - (Stars:) supernovae: general - Equation of state - Nuclear reactions, nucleosynthesis, abundances - Methods: numerical
@@ 1 Introduction
Building stellar evolution models to reproduce observations remains a difficult task, due to a relatively large amount of free parameters and approximations used in the modelling. An increase in the quality and quantity of observations of stellar objects, (e.g. the increasing catalogue of observations by the GAIA collaboration, Gaia Collaboration et al. 2016, 2018, 2021), refines the global constraints that stellar models must satisfy (most of which refer to photospheric conditions), as can be found in recent studies (Dal Tio et al. 2022; Massey et al. 2023; Yang et al. 2019). The development of asteroseismology techniques, reviewed by Aerts et al. (2019), can provide an insight into the internal characteristics of stars that also helps constrain stellar modelling (e.g. Burssens et al. 2023; Bowman 2023; Eggenberger et al. 2019a, b). Furthermore, the observations at ever larger redshifts provided by the JWST give hints about the early stellar generations in the universe that may have produced, for example, the seeds of supermassive black holes or strongly nitrogen enriched regions (see for instance Ji et al. 2024; Jeon et al. 2024; Nandal, Devesh et al. 2024; Nandal et al. 2024). Among the most prominent tasks for stellar models is to verify the constraints provided from all of these observations.
@@
One major difficulty for massive stars above 8 $M_{\odot}$ is that the stellar core in the final phases of evolution, (i.e. from the beginning of carbon fusion in the core), is obscured from view deep inside the large stellar envelopes. The impact of mixing and nuclear burning in the stellar interior therefore cannot be well constrained by observations. Understanding the evolution of the stellar core in this regime heavily relies on numerical models, whose validation chiefly depends on observations of the subsequent evolution. Specifically, pre-supernova models are further employed as the starting points to predict the fate after core bounce and through to the supernova explosions (e.g. Ugliano et al. 2012; Sukhbold & Woosley 2014; Sukhbold et al. 2016, 2018) as well as the multi-messenger signal of the post-explosion evolution (e.g Murphy et al. 2009; Yakunin et al. 2010; Scheidegger et al. 2010; Mueller et al. 2013; Lund et al. 2012; Cerdá-Durán et al. 2013; Tamborra et al. 2013, 2014b, 2014a; Kuroda et al. 2016; Mirizzi et al. 2016; Kashiyama et al. 2016; Janka et al. 2016; Andresen et al. 2017; Kuroda et al. 2017; Janka 2016; Kuroda et al. 2017; Pan et al. 2018; O’Connor & Couch 2018; Powell & Müller 2019; Radice et al. 2019; Morozova et al. 2018; Walk et al. 2018; Andresen et al. 2019; Lin et al. 2020; Walk et al. 2019; Müller 2019; Andersen et al. 2021; Powell & Müller 2020; Pan et al. 2021; Mezzacappa et al. 2020; Schneider et al. 2021; Andersen et al. 2021; Jardine et al. 2022; Mezzacappa et al. 2023; Bugli et al. 2023, 2023; Vartanyan et al. 2023). With these predictions, we may compare observations of the observed statistics of neutron stars and black holes, including constraints on their spins, supernova explosion properties, predicted gravitational waveforms, neutrinos or nucleosynthetic yields with the obtained results, and hence constrain the pre-supernova models (e.g. Smartt et al. 2009; Smartt 2015; Radice et al. 2019; Fonseca et al. 2023; Abe et al. 2024; Reichert et al. 2021, 2023, 2024; Sieverding et al. 2023).
@@
Late-phase stellar evolution models have been studied by numerous authors using a variety of different stellar evolution codes. With GENEC, Hirschi et al. (2004) and Ekström et al. (2012), pushed models beyond silicon-core burning. For the KEPLER stellar evolution code studies such as Heger et al. (2001a),Sukhbold & Woosley (2014), and Sukhbold et al. (2018) have produced a large number of pre-supernova models. The development of MESA in Paxton et al. (2011, 2013, 2015, 2018, 2019) has also produced grids of pre-supernova models. All these studies have provided initial models to probe the subsequent dynamical evolution of collapsing massive stars.
@@
The properties of progenitor stars at the moment of collapse are crucial in order to determine the outcome of the post-collapse evolution of massive stars. Among these properties we mention, the density, the rotational and magnetic profiles, the mass of the iron core and its radius (ordinarily quantified in terms of its compactness parameter O’Connor & Ott 2011), or its electron fraction profile (e.g. Baron & Cooperstein 1990, among others). Supernova explosions are very sensitive to the progenitor structure; the entropy, electron fraction, equation of state, and neutrino physics all have a strong impact (Cooperstein & Wambach 1984; Baron et al. 1985; Bethe 1990). The stellar profiles of these quantities result from a complex balance of nuclear burning stages and the creation, spatial dimension, and temporal sequence of different convective layers (some of which may mutually interact). Convection in one-dimensional stellar evolution codes is implemented assembling together the Schwarzchild-Ledoux (local) instability criterion (either the Schwarzshild or the Ledoux criterion is used, but not together), and basic prescriptions for diffusion of chemical species and angular momentum (in addition to parametric models for semi-convection, overshooting). Since the incorporation of these elements varies from one stellar evolution code to another, more often than desirable, the pre-supernova state for seemingly similar zero age main sequence (ZAMS) conditions differs among different groups (Jones et al. 2015; Chieffi & Limongi 2020).
@@
In this paper, we show the details of the implementation of an extended equation of state (EoS) in GENEC covering the whole density-temperature plane required to reach core collapse for massive stars. This EoS patches together two pre-existing EoSs, which include complementary physical aspects, along with some common elements (Sect. 2). We show that while the local values of the EoS recovered thermodynamic variables do not show large discrepancies, their derivatives may have significant deviations, specifically derivatives impacting the criteria used to distinguish among zones that are unstable or stable to convection. Additionally, we extend the coverage of the radiative and conductive opacities to high-temperatures and densities to include the effects of degeneracy and high-energy physics, which play an important role in the final phases of stellar life (Sect. 3). With the aim of approximately covering the most advanced burning phases of the evolution of massive stars, we also extend the existing reduced nuclear reaction network to include electron capture on heavy nuclei (Sect. 4). Therefore, the nuclear network approximately accounts for the reduction of the electron fraction ($Y_{e}$) and consequently the electron pressure, which induces core collapse. To validate these improvements, using GENEC, we evolved several models of various massive stars until the brink of core collapse, which we compare to corresponding models evolved with MESA (Sect. 5). We discuss our results and present our conclusions in Sect. 6. Since the implementation details of the previous default EoS in GENEC are not easily accessible in electronic form, we highlight the most relevant features of that EoS in the Appendix B.
@@
2 A new EoS to cover the regime from partial ionisation to relativistic electron degeneration
@@
Stellar matter is usually modelled as a mixture of ideal (baryonic) gases, radiation, and electrons/positrons. It encompasses regimes of partial to full ionisation. The current version of the GENEC stellar evolution code employs, by default, the Dichte [^1] The name Dichte comes from the German word for density as the code was developed within the group of Kippenhahn and the EoS returns the density when one inputs pressure, temperature, and composition. EoS (D-EoS, hereafter), which accounts for the partial ionisation regime of certain light elements, (hydrogen, helium, carbon, oxygen, nitrogen, and magnesium) in the stellar envelope (see Appendix B for details), as well as fully ionised medium of the stellar interior. We note that extending the treatment of partial ionisation for heavier elements would be welcome, as discussed in Jermyn et al. (2021), but this is a matter for future work. Two important regimes for massive star evolution are not covered by the D-EoS, namely the electron degeneracy in the high-density (i.e. relativistic) regime and the effects of electron-positron pair creation in the very high-temperature regime. To cover this region of the density-temperature parameter space, we have extended the existing GENEC EoS by matching it with the Timmes EoS (T-EoS, hereafter) (Timmes et al. 2000) in the deep stellar interiors. The latter EoS computes the electron-positron contribution based upon an interpolation table of the Helmholtz free energy (ensuring thermodynamic consistency). The T-EOS also includes the Coulomb energy term, absent from the D-EoS, which can be relevant in high-density and low-temperature conditions, however, such conditions are generally not reached in the models presented in this work.  [^2] The implementation of the T-EOS in GENEC, employs the subroutines read_helm_table, read_helm_iontable, and helmeos from the freely available version on the web at https://cococubed.com/codes/eos/helmholtz.tbz, version of 2022-01-08. It should be noted that the Timmes EoS (T-EoS hereafter) assumes total ionisation of the plasma, which is not the case in the whole stellar interior. To allow for the EoS to deal with the partial ionisation in the stellar envelope, we employ a switch between the two EoSs. According to Jermyn et al. (2021) the ionisation effects become relevant for temperatures $T<T^{\text{ion}}\approx 10^{4}Z_{j}^{2}\,$ and densities $\rho<\rho^{\text{ion}}\approx 3A_{j}Z_{j}^{3}\,\text{g\,cm}^{-3}$, where $Z_{j}$ and $A_{j}$ are the atomic number and atomic mass of the considered element. Using these criteria on the ionisation conditions as approximate guidelines, and trying to minimise the jump between the two different EoSs, if the following condition is verified, we use the D-EoS
$\hskip 65.44142pt\rm log(\rho)\leq 2.8\ \ \rm{or}\ \ \rm log(T)\leq 7.55,$ (1)
@@
Thus, noting that the switch to full ionisation is handled according to Eq. (43) in the D-EoS, the partial ionisation regime for hydrogen and helium is entirely addressed by this EoS (and not by the T-EoS). In the following, we collectively refer to the coupled EoS employed in GENEC as the DT-EoS.
@@ 2.1 Electronic contribution to the pressure
The T-EoS represents an improvement over the D-EoS when it comes to computing the electronic pressure (see Appendix B). While the D-EoS employs analytic estimations for the evaluation of the Fermi-Dirac integrals that are used to compute the electron contribution to the thermodynamic variables, the T-EoS uses a more accurate table for the leptonic contribution including electrons and positrons.
In Fig. 1 we show the variations of the electron-positron pressure contribution as a function of $\rho$, and $T$, between the two EoSs. The part of the $(\rho,T)$-plane represented in that figure falls fully inside the region where we employ the T-EoS. For the computation of the pressure we take the case of pure $\rm{}^{16}O$, for which $\overline{A}=16$ and $\overline{Z}=8$, although changing the composition does not alter the conclusions made here. The quantity represented in Fig. 1 is
$\chi=\frac{P_{\rm e}^{T}}{P^{T}}\times\frac{\left|P_{\rm e}^{D}-P_{\rm e}^{T}\right|}{P_{\rm e}^{T}},$ (4)
@@
[…]
equal to unity. Below this line Coulomb corrections contribute significantly to the EoS (see, e.g. Maeder 2009, Eq. (7.120) for further details).
@@ 2.2 Impact on density
The change of EoS also impacts the density profile in the star, a crucial element for the determination of the stellar structure. The T-EoS takes as input a pair of values $(\rho,T)$ to construct the Helmholtz free energy and evaluate the pressure $P^{\rm helm}$ (Timmes & Swesty 2000). On the other hand, the D-EoS takes as input the pair of values $(P,T)$ and iterates over the degeneracy parameter, $\Psi$, or electron chemical potential, [^3] We recall that the degeneracy parameter is the electron chemical potential in units of $k_{\mathrm{B}}T$, where $k_{\mathrm{B}}$ is the Boltzmann constant. to find the density ($\rho^{\rm D}$) that satisfies the equations shown in the Appendix B. In Fig. 2 we show the logarithm of the difference $\rho^{\rm D}/\rho^{\rm T}-1$. Where $\rho^{\rm T}$ is the input density for the T-EoS and $\rho^{\rm D}$ the output density of the D-EoS. $\rho^{\rm D}$ is obtained by using as input value $P^{\rm T}$ (i.e. the pressure computed using the T-EoS for the thermodynamic conditions $(\rho^{\rm T},T)$). We note that the temperature, $T$, has been fixed in all the process of comparison.
@@
We see in Fig. 2 that in a large part of the $(\rho,T)$ plane $\log(\rho^{\rm D}/\rho^{\rm T}-1)<-2$. It is only in the upper left region, where the pressure is dominated mainly by the radiation contribution, that the D-EoS begins to diverge strongly from the thermodynamic consistent T-EoS. In this region, the contributions made by the positrons to the pressure can be large. In general, the presence of positrons in the EoS will affect the electron chemical potential and thus the degeneracy parameter. The lack of positrons in the D-EoS, and the analytic expressions used to evaluate the Fermi-Dirac integrals (Appendix B), leads to large differences in the estimation of the degeneracy parameter between the two EoSs. For example, at the point $(\rho,T)=(10^{5}\ \rm g\,cm^{-3},10^{9.5}\ \rm K)$, for a pure silicon composition, the T-EoS returns $\psi^{T}=-1.8$ whereas the D-EoS returns $\psi^{D}=-1.2$. This degeneracy parameter with the D-EoS gives a density of roughly $6\times 10^{5}\,\rm g\,cm^{-3}$ a considerable change to the Timmes input of $1\times 10^{5}\,\rm g\,cm^{-3}.$ Thus using the Timmes EoS (or more sophisticated alternatives; see e.g. Jermyn et al. 2021) in this region is mandatory to obtain a correct evaluation of the density and, thereby, the correct stellar structure.
@@
Most of the stellar interior of our models are not affected by this region of strong differences in the density. However, in Fig. 2 we super-impose the $(\rho,T)$ internal structure of a $15M_{\odot}$ model computed with the D-EoS in magenta and a model using the DT-EoS. The main shift between these models comes from the location of the interfaces, represented by sharp variations in the $(\rho,T)$ curve. The size of jump at these interfaces in density and temperature is also different between the two models using different EoSs. The properties at shell interfaces are as we see in Sect. 6 crucial for understanding supernova explosions of these progenitors.